Systems of Linear Equations : Solve Algebraically
2x^2+y^2=17
x^2+2y^2=22

Answer by checkley79(3341)

2x^2+y^2=17 multiply by -2 & add.
x^2+2y^2=22
-4x^2-2y^2=-34
-------------------
-3x^2=-12
x^2=-12/-3
x^2=4
x=sqrt4
x=2 ans.
4+2y^2=22
2y^2=22-4
2y^2=18
y^2=18/2
y^2=9
y=sqrt9
y=3 ans.
Proof:
2*4+9=17
8+9=17
17=17